OpenStudy (anonymous):

At a manufacturing plant, parts are regularly tested. Past records indicate that two percent of all parts tested are found to be defective. In a test batch of four hundred parts, how many are expected to be defective?

OpenStudy (anonymous):

2 percent means 2 out of a hundred. so two percent of x will be: \[x \times \frac{ 2 }{ 100 }.\] 2% of 400 will be: \[400 \times \frac{ 2 }{ 100 } = 8.\]